# Mesoscopic superfluid to superconductor transition

### v1

@@ I Introduction
Minimal model for SF.– The BH model serves as a minimal model for the explanation of SF. Namely, the superflow becomes metastable thanks to the inter-particle interactions. But if the interaction is further increased, the ground state changes from coherent condensation, to a fragmented site-occupation, which is the SF-MI transition [11, 12, 13]. Formally speaking, this transition is abrupt only for an infinite chain, but its mesoscopic version is clearly apparent even for the two-site model, as pointed out long ago by Leggett [1], who called it a transition from Josephson-regime to Fock-regime [14, 15]. For a ring with several sites, condensation in an excited momentum orbital implies the feasibility of superflow. Additionally, one has to address the significance of the underlying chaos. Recent studies of the chaos-mediated SF-MI transition are [16, 17] and [18]. The latter adopts a tomographic approach that we would like to pursue.
@@
Minimal model for SC.– The question arises, what is the minimal model for the discussion of the SF-SC-MI transition and the Meissner effect. We take here literally the old idea that SC can be regarded as arising from condensation of charged bosons. Accordingly, the minimal model consists of a BH ring coupled to a single EM mode of a cavity. The latter is modeled as an LC oscillator. This configuration is enough to demonstrate the emergence of SC. The dimensionless parameter that characterizes the coupling between the BH ring the the EM mode is a dimensional fine structure constant $\alpha$. This parameter controls the crossover from the SF regime to an SC regime. The excitation frequencies of the system are affected by this EM interaction, which constitutes a mesoscopic version of the Anderson-Higgs mechanism. On top we have the inter-particle interaction that can induce a transition to an MI regime.
@@
Quantum chaos.– The SF-MI transition is apparent also in the parametric evolution of the excited states. In this context, quantum chaos is an aspect that should not be overlooked [19, 20, 21, 22, 23, 24, 25, 26, 16, 17]. The common “quantum chaos” practice to characterize a system is to look on its energy level statistics. But such approach has two issues: (i) It provides numerically clear results only for rather simple systems with large number ($N$) of particles, such that the spectrum in the range of interest is dense enough; (ii) It does not allow classification of the eigenstates in the typical situation of underlying mixed chaotic and quasi-regular dynamics. In the present context, one would like to highlight meta-stable SF states that are embedded in a background of FR states.
@@
Tomography.– As opposed to the common practice, our approach is numerically cheap and flexible enough to deal with small complex systems. We use the term quantum phase-space tomography [18] in order to emphasize that the inspection of the spectrum is not limited to “level statistics”, neither focused on the ground state, and possibly reveals the relation to the underlying phase space structure. The latter serves as a classical skeleton for the quantum eigenstates.
@@
Outline.– The circuit Hamiltonian and its parameters are introduced in Sec. (II) and Sec. (III). The big picture regarding SF/SC superflow is provided in Sec. (IV). The topography of the energy landscape is discussed in Sec. (V), with focus on the Landau perspective. The semiclassical SF-SC border is explained in Sec. (VI). Tomography measures are defined in Sec. (VII) and applied in Sec. (VIII). The various regions in the quantum spectrum are further elaborated in Sec. (IX). The mesoscopic Meissener effect and its relation to the Anderson-Higgs mechanism are discussed in Sec. (X). We finalize with a summary in Sec. (XI). The extra appendices Appendix A and Appendix B and Appendix C help to place this work in the common context of SC studies.
@@ V The energy landscape
The energy landscape is dictated by 3 relevant energy scales
@@ V.2 Valley structure
The potential floor feature a valley structure. There is a major valley around the $k_{0}$ condensate at $\phi_{0}$, and possibly there are several metastable valleys around the $k_{m}$ condensates at $\phi\sim\phi_{m}$}. An illustration is given in Fig. 1a where the lowest valley is indicated by a green color, and the two metastable valleys are indicated by red and blue colors. The energies of the meta-stable minima are $E\sim W_{\Phi}m^{2}$}, and they are separated by barriers whose height is of order $W_{K}$. Accordingly, the energy floor features $\sim w_{L}$ major valleys. For large enough $\alpha$ there are $L$ major valleys, each correspond to a different condensate. They are located at the smallest values of $\phi_{m}$}. Due to the discreteness of the ring there might be also secondary valleys, for larger |m|{|m|} values.
@@ V.3 Landau Grooves
The major washboard-sections that correspond to $k_{m}$ condensates are separated by intermediate washboard-sections that correspond to fragmented occupation. The 3rd term of Eq. (8) implies that due to an energy cost $W_{U}$, a condensates might have lower energy compared with fragmented states of the nearby washboard-sections. To be more precise, the valley that is associated with the $k_{m}$ condensate looks like a groove that is formed in the potential floor $V(n_{k},\phi)$}. This groove is centered at $\phi_{m}$, and its stretch along the $\phi$ direction is determined by the Landau criterion as explained below.
@@
Consistency with the large $L$ approximation is easily verified. Optionally, given $L$, the critical value $\phi_{c}$} increases with $u_{L}$ from the minimal value $\phi_{c}\sim\pi$} to the limiting value $\phi_{c}\sim(L/2)\pi$}.
@@ VII.2 Condensation measures
Roughly, $1/\mathcal{S}_{purity}$ is the number of participating orbitals, such that $\mathcal{S}_{purity}=1$ implies condensation in a single orbital. Irrespective of that $\mathcal{S}_{ent}=1$ implies that the ring-osc eigenstate can be factorized, which is the case for either full SF condensation or strict MI fragmentation. We define $\mathcal{N}_{ent}=1/\mathcal{S}_{ent}$}, and note that $\log[\mathcal{N}_{ent}]$} is like entanglement entropy.
@@ VIII Tomography of the spectrum
We demonstrate numerically spectrum tomography for the simplest circuit, namely, a trimer ring (L=3L{=}3). Depending on parameters such ring can support either an SF ground-state or SC multi-stability. In practice $\alpha$ of Eq. (6) is very small, but the SF-SC transition is feasible because $N\alpha$ can be large. In the numerics $N$ is limitted, and therefore we artificially compensate it by setting unrealistic large $\alpha$. Nevertheless, the outcome is semiclassically equivalent, and therefore satisfactory for the purpose of illustration.
@@
Figure 2: The energy landscape: We consider u=0u{=}0 trimer with N=12{N=12} particles coupled to an EM mode. The other parameters are $\omega_{0}=0.48$} and $\alpha=2$}, while $N_{osc}=150$}. The upper panels provide images for eigenstates of $\mathcal{H}_{osc}$ Eq. (14) assuming that the particles are condensed in one of the three $k$ orbitals. Each row is the $\phi$ probability distribution. The rows are ordered by energy. The lower left panel combines the 3 upper upper panels. It would be the full spectrum if we had a single particle. For N=12{N=12} one has to combine the spectra that are associated with all possible occupations. The outcome (zoomed) is the middle lower panel. In the right lower panel we display the result for $\alpha=1/2$}, where SC multi-stability is absent.
Figure 3: Tomography of the spectrum: The spectrum for trimer with $N=12$ particles, coupled to an EM mode with $\omega_{0}=0.48$ and $N_{osc}=150$. Each eigenstate is represented by a point that is positioned according to its energy index $\nu$ and its $\left\langle\phi\right\rangle$. The RGB color-code indicates the occupation of the orbitals. In the first row the interaction is u=1/2{u=1/2} and the coupling is $\alpha=2$}. The two right panels are zoomed versions of the left panel, where the upper false eigenstates that appear due to finite $N_{osc}$ are truncated, and the SC multi-stability is highlighted. In the second row the interaction is u=3{u=3} and the couplings, from left to right, are $\alpha=1/2,2,20$}. The panels are zoomed over the first 30003000 states. This set demonstrates the crossover from SF to SC multi-stability. Note that the range of the horizontal axis has been adjusted.
@@
Figure 4: The formation of SC multi-stability: (a) Each column is an image of the spectrum for a different value of $\alpha$. The energy levels (indexed by $\nu$) are RGB color-coded, reflecting orbital occupation as explained in the text. The interaction between the particles is weak (u=0.01{u=0.01}). The other parameters are $\omega_{0}=0.48$}, and N=12{N=12} and $N_{osc}=60$}. The states are in the range $0<\nu<0.05\mathcal{N}$}, namely, the 277 lowest levels out of 5551 are displayed. (b) The dependence of the spectrum on $u$ for representative values of EM coupling, namely, $\alpha=0.1,2,20$}. This set demonstrates the crossover from SF to SC, and how SF/SC is diminished due to $u$. (c) The information of panel b3 is displayed in a different way. Each point is positioned according to the normalized orbital occupations, and the color-code indicates $u$. As the interaction is increased multi-stability is diminished.
Figure 5: Multi-stability versus interaction. Optional diagrams that show the $u$ dependence of the spectrum. The 3 panels are: (a) The normalized occupation $n_{0}$ of the zero momentum orbital; (b) The purity $S_{purity}$ serves as a condensation measure; (c) The EM mode flux $\left\langle\phi\right\rangle$ serves to identify SC symmetry breaking. The parameters are N=12{N=12}, and $N_{osc}=60$}, and $\alpha=2$}, and $\omega_{0}=0.48$}.
@@ VIII.1 Spectra
Fig. 3 displays images for selected eigenstates of an $L=3$ circuit for $U=0$. Namely, each row is an image of the $\phi$ probability distribution for an eigenstate of Eq. (14). The zoomed panels show images of all the low-lying eigenstates for $\alpha=1/2$} and $\alpha=2$}. The latter exhibits SC multi-stability as opposed to the former that feature an SF ground-state. A numerical remark is in order here. Mathematically speaking quasi-degenerated SC metastable states should combine into cat-state superpositions that feature $\left\langle\phi\right\rangle=0$}, but we care to introduce a numerically negligible bias $\Phi_{ext}$ that allows ‘spontaneous’ breaking of this symmetry.
Fig. 3 displays the full spectrum for non-zero interaction. Each eigenstate is represented by a single point that is RGB color-coded such that red, green, and blue indicate the respective occupations $(n_{-},n_{0},n_{+})$}. The $n_{k}$ are not “good quantum numbers” but rather are expectation values. The selected set of panels clearly demonstrates the crossover from SF to SC multi-stability.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 6: Ergodicity and Entanglement: Left panels from top to botom: images of $\overline{\mathcal{M}}$ and $\mathcal{M}$ and $\mathcal{N}_{ent}$. The parameters are the same as in Fig. 5. Right panels: scatter diagrams color-coded by $u$, for inspecting correlations between the 3 measures. The upper diagram allows to characterize the degree of ergodicity. For reference we plot a diagonal magenta line that indicates full ergodicity. The two other diagrams demonstrate that $\mathcal{N}_{ent}$ is more correlated with $\overline{\mathcal{M}}$ than with $\mathcal{M}$.
@@ VIII.2 Regimes
Fig. 5b demonstrates how the multi-stability is diminished as $u$ is increased. This is the transition to the MI regime where the eigenstates get fragmented in the Fock site basis. Optional visualizations of this dependence in provided in Fig. 5 where we inspect separately the occupation $n_{0}$ of the zero momentum orbital; the purity $S_{purity}$ that serves as a condensation measure; and $\left\langle\phi\right\rangle$ that serves to identify the SC symmetry breaking.
@@ VIII.3 Ergodicity
The underlying phase space is in general chaotic and possibly mixed with quasi-regular motion. In order to diagnose this aspect we provide in Fig. 6 the (u,E){(u,E)} diagrams for both $\overline{\mathcal{M}}$ and $\mathcal{M}$. Then we construct a scatter diagram to characterize the degree of ergodicity. We see that accessible phase space region expands at the transition, indicated by $\overline{\mathcal{M}}$. Crossing the diagram in the increasing $u$ direction, the ergodicity increases slowly as indicated by the ratio $\mathcal{M}/\overline{\mathcal{M}}$.

### v2

@@ I Introduction
The BH model serves as a minimal model for the explanation of SF. Namely, the superflow becomes metastable thanks to the inter-particle interactions. But if the interaction is further increased, the ground state changes from coherent condensation, to a fragmented site-occupation, which is the SF-MI transition [11, 12, 13]. Formally speaking, this transition is abrupt only for an infinite chain, but its mesoscopic version is clearly apparent even for the two-site model, as pointed out long ago by Leggett [1], who called it a transition from Josephson-regime to Fock-regime [14, 15]. For a ring with several sites, condensation in an excited momentum orbital implies the feasibility of superflow. Additionally, one has to address the significance of the underlying chaos. Recent studies of the chaos-mediated SF-MI transition are [16, 17] and [18]. The latter adopts a tomographic approach that we would like to pursue.
Minimal model for SC.–
@@
The question arises, what is the minimal model for the discussion of the SF-SC-MI transition and the Meissner effect. We take here literally the old idea that SC can be regarded as arising from condensation of charged bosons. Accordingly, the minimal model consists of a BH ring coupled to a single EM mode of a cavity. The latter is modeled as an LC circuit. This configuration is enough to demonstrate the emergence of SC. The dimensionless parameter that characterizes the coupling between the BH ring the the EM mode is a dimensional fine structure constant $\alpha$. This parameter controls the crossover from the SF regime to an SC regime. The excitation frequencies of the system are affected by this EM interaction, which constitutes a mesoscopic version of the Anderson-Higgs mechanism. On top we have the inter-particle interaction that can induce a transition to an MI regime.
Quantum chaos.–
@@
The SF-MI transition is apparent also in the parametric evolution of the excited states. In this context, quantum chaos is an aspect that should not be overlooked [19, 20, 21, 22, 23, 24, 25, 26, 16, 17]. The common “quantum chaos” practice to characterize a system is to look on its energy level statistics. But such approach has two issues: (i) It provides numerically clear results only for rather simple systems with large number ($N$) of particles, such that the spectrum in the range of interest is dense enough; (ii) It does not allow classification of the eigenstates in the typical situation of underlying mixed chaotic and quasi-regular dynamics. In the present context, one would like to highlight meta-stable SF states that are embedded in a background of FR states.
Tomography.–
@@
As opposed to the common practice, our approach is numerically cheap and flexible enough to deal with small complex systems. We use the term quantum phase-space tomography [18] in order to emphasize that the inspection of the spectrum is not limited to “level statistics”, neither focused on the ground state, and possibly reveals the relation to the underlying phase space structure. The latter serves as a classical skeleton for the quantum eigenstates.
Outline.–
@@
The circuit Hamiltonian and its parameters are introduced in Sec. (II) and Sec. (III). The big picture regarding SF/SC superflow is provided in Sec. (IV). The topography of the energy landscape is discussed in Sec. (V), with focus on the Landau perspective. The semiclassical SF-SC border is explained in Sec. (VI). Tomography measures are defined in Sec. (VII) and applied in Sec. (VIII). The various regions in the quantum spectrum are further elaborated in Sec. (IX). The mesoscopic Meissener effect and its relation to the Anderson-Higgs mechanism are discussed in Sec. (X). We finalize with a summary in Sec. (XI). The extra appendices Appendix A and Appendix B and Appendix C help to place this work in the common context of SC studies.
@@ V The energy landscape
The energy landscape is dictated by 3 energy scales
@@ V.1 Potential floor
The Born-Oppenheimer washboard sections are illustrated in Fig. 1 for an $L=3$ ring. The RGB color-code reflects the occupation such that red, green and blue indicate condensation in one of the 3 orbitals.
@@ V.2 Valley structure
The potential floor feature a valley structure. There is a major valley around the $k_{0}$ condensate at $\phi_{0}$, and possibly there are several metastable valleys around the $k_{m}$ condensates at $\phi\sim\phi_{m}$}. In the illustration of Fig. 1a the lowest valley is formed by the green colored sections, and the two metastable valleys are formed by the red and blue colored sections. The energies of the meta-stable minima are $E\sim W_{\Phi}m^{2}$}, and they are separated by barriers whose height is of order $W_{K}$. Accordingly, the energy floor features $\sim w_{L}$ major valleys. For large enough $\alpha$ there are $L$ major valleys, each correspond to a different condensate. They are located at the smallest values of $\phi_{m}$}. Due to the discreteness of the ring there might be also secondary valleys, for larger |m|{|m|} values.
In the numerical demonstration we forcus on the 3 major vallies of an $L=3$ ring. If we used larger $L$, there would be more ‘color-distinct’ major valleys. Other than that, as long as the ‘transitions’ in Eq. (8) are neglected, the overall topography of potential floor depends on $L$ merely via the classical dimensionless parameters. If we take the ‘transitions’ into account, then $L$ has a secondary effect on the appearance of the Landau grooves, as further discussed below. These grooves are important for the understanding of SF metastability, but less important for the discussion of SC.
@@ V.3 Landau Grooves
The major washboard-sections that correspond to $k_{m}$ condensates are separated by intermediate washboard-sections that correspond to fragmented occupation. The 3rd term of Eq. (8) implies that due to an energy cost $W_{U}$, a condensates might have lower energy compared with fragmented states of the nearby washboard-sections. To be more precise, the valley that is associated with the $k_{m}$ condensate is centered at $\phi_{m}$, and its stretch along the $\phi$ direction is determined by the Landau criterion as explained below. The stretch looks like a groove along the potential floor $V(n_{k},\phi)$}.
@@
Consistency with the large $L$ approximation is easily verified. On the other hand, given $L$, the critical value $\phi_{c}$} increases with $u_{L}$ from the minimal value $\phi_{c}\sim\pi$} to the limiting value $\phi_{c}\sim(L/2)\pi$}.
@@ VII.2 Condensation measures
Roughly, $1/\mathcal{S}_{purity}$ is the number of participating orbitals, such that $\mathcal{S}_{purity}=1$ implies condensation in a single orbital. Irrespective of that, $\mathcal{S}_{ent}=1$ implies that the ring-osc eigenstate can be factorized, which is the case for either full SF condensation or strict MI fragmentation. We define $\mathcal{N}_{ent}=1/\mathcal{S}_{ent}$}, and note that $\log[\mathcal{N}_{ent}]$} is like entanglement entropy.
@@ VIII Tomography of the spectrum
We demonstrate numerically spectrum tomography for the simplest circuit, namely, a trimer ring (L=3L{=}3). Depending on parameters such ring can support either an SF ground-state or SC multi-stability. If we take Eq. (6) literally, the realistic value of $\alpha$ is very small. Considering an $\mathcal{L}\sim 10\mu\text{m}$} ring it comes out $\alpha\sim 10^{-4}$} for an optical cavity, and much smaller for Microwave cavity. In a numerical demonstration one is limited by $N\sim 50$}, and therefore, to illustrate an SF-SC transition, we artificially compensate by setting unrealistic large $\alpha$. In any case, one should be aware that to regard the model as an actual representation of a realistic experimental system is rather challenging.
@@
Figure 2: The energy landscape: We consider u=0u{=}0 trimer with N=12{N=12} particles coupled to an EM mode. The other parameters are $\omega_{0}=0.48$} and $\alpha=2$}, while $N_{osc}=150$}. The upper panels provide images for eigenstates of $\mathcal{H}_{osc}$ Eq. (14) assuming that the particles are condensed in one of the three $k$ orbitals. Each row of a panel is an image of the $\phi$ probability distribution. The rows within each panel are ordered by energy. The lower left panel combines the 3 upper upper panels. It would be the full spectrum if we had a single particle. For N=12{N=12} one has to combine the spectra that are associated with all possible occupations. The outcome (zoomed) is the middle lower panel. In the right lower panel we display the result for $\alpha=1/2$}, where SC multi-stability is absent.
Figure 3: Tomography of the spectrum: The spectrum for trimer with $N=12$ particles, coupled to an EM mode with $\omega_{0}=0.48$ and $N_{osc}=150$. Each eigenstate is represented by a point that is positioned according to its energy index $\nu$ and its $\left\langle\phi\right\rangle$. The RGB color-code indicates the occupation of the orbitals. In the first row the interaction is u=1/2{u=1/2} and the coupling is $\alpha=2$}. In the left-most panel the arrows are pointing towards the green ground-state and towards the major red and blue meta-stable states. The two right panels are zoomed versions of the left panel, showing with better resolution this SC multi-stability, and excluding the upper false eigenstates that appear due to finite $N_{osc}$. In the second row the interaction is u=3{u=3} and the couplings, from left to right, are $\alpha=1/2,2,20$}. The panels are zoomed over the first 30003000 states. This set demonstrates the crossover from SF to SC multi-stability. Note that the range of the horizontal axis has been adjusted.
@@
Figure 4: The formation of SC multi-stability: (a) Each column of the panel is an image of the spectrum for a different value of $\alpha$. The energy levels (indexed by $\nu$) are RGB color-coded, reflecting orbital occupation as explained in the text. The interaction between the particles is weak (u=0.01{u=0.01}). The other parameters are $\omega_{0}=0.48$}, and N=12{N=12} and $N_{osc}=60$}. The states are in the range $0<\nu<0.05\mathcal{N}$}, namely, the 277 lowest levels out of 5551 are displayed. (b) The dependence of the spectrum on $u$ for representative values of EM coupling, namely, $\alpha=0.1,2,20$}. This set demonstrates the crossover from SF to SC, and how SF/SC is diminished due to $u$. (c) The information of panel b3 is displayed in a different way. Each point is positioned according to the normalized orbital occupations, and the color-code indicates $u$. As the interaction is increased multi-stability is diminished.
Figure 5: Multi-stability versus interaction. Additional diagrams that show the $u$ dependence of the spectrum. The 3 panels are: (a) The normalized occupation $n_{0}$ of the zero momentum orbital; (b) The purity $S_{purity}$ serves as a condensation measure; (c) The EM mode flux $\left\langle\phi\right\rangle$ serves to identify SC symmetry breaking. The parameters are N=12{N=12}, and $N_{osc}=60$}, and $\alpha=2$}, and $\omega_{0}=0.48$}.
Figure 6: Ergodicity: From left to right: Images of $\overline{\mathcal{M}}$ and $\mathcal{M}$. The parameters are the same as in Fig. 5. Scatter diagrams color-coded by $u$, for inspecting the correlation between the 2 measures. For reference we plot a diagonal magenta line that indicates full ergodicity. The trend in the “quantum chaos regime” is indicated by an arrow that points in the direction of increasing $u$.
Figure 7: Entanglement: From left to right: Image of $\mathcal{N}_{ent}$. The parameters are the same as in Fig. 5. The other two panels are scatter diagrams color-coded by $u$, for demonstrating that $\mathcal{N}_{ent}$ is more correlated with $\overline{\mathcal{M}}$ than with $\mathcal{M}$.
@@ VIII.1 Spectra
Fig. 3 displays images for selected eigenstates of an $L=3$ circuit for $U=0$. Namely, each row is an image of the $\phi$ probability distribution for an eigenstate of Eq. (14). The upper eigenstates are false: they appear due to the finite truncation $N_{osc}$. Namely, due to this truncation the spectrum is bounded from above. The zoomed panels exclude the false eigenstates. They show images of all the low-lying eigenstates for $\alpha=1/2$} and $\alpha=2$}. The latter exhibits SC multi-stability as opposed to the former that feature an SF ground-state. A numerical remark is in order here. Mathematically speaking quasi-degenerated SC metastable states should combine into cat-state superpositions that feature $\left\langle\phi\right\rangle=0$}, but we care to introduce a numerically negligible bias $\Phi_{ext}$ that allows ‘spontaneous’ breaking of this symmetry.
Fig. 3 displays the full spectrum for non-zero interaction. Each eigenstate is represented by a single point that is RGB color-coded such that red, green, and blue indicate the respective occupations $(n_{-},n_{0},n_{+})$}. The $n_{k}$ are not “good quantum numbers” but rather are expectation values. The selected set of panels clearly demonstrates the crossover from SF to SC multi-stability. In all subsequent figures we care to exclude the upper false eigenstates that appear due to the finite truncation $N_{osc}$. Whatever is displayed is independent of this truncation, and the identification of the various regimes (SF-SC-FR-MI) is not affected.
@@ VIII.2 Regimes
Fig. 5b demonstrates how the multi-stability is diminished as $u$ is increased. This is the transition to the MI regime where the eigenstates get fragmented in the Fock site basis. Further visualizations of this dependence in provided in Fig. 5 where we inspect separately the occupation $n_{0}$ of the zero momentum orbital; the purity $S_{purity}$ that serves as a condensation measure; and $\left\langle\phi\right\rangle$ that serves to identify the SC symmetry breaking.
@@ VIII.3 Ergodicity
The Hamiltonian is integrable in both limits of very small and very large $u$. In particular, in the large $u$ regime, where the site-occupations become good quantum numbers, one encounters the transition to MI. In Ref[18] we have provided a very detailed tomographic account of this transition, trying to identify eigenstates that reflect particle-hole excitations, and also amplitude and phase modes that appear at the entrance to the MI regime. It seems that for charged bosons the emerging picture is similar, we therefore do not dwell further on this theme. For intermediate values of $u$ the underlying phase space is chaotic and possibly mixed with quasi-regular motion. In order to diagnose this aspect we provide in Fig. 7 the (u,E){(u,E)} diagrams for both $\overline{\mathcal{M}}$ and $\mathcal{M}$. We see that the accessible phase space region expands for intermediate values of $u$, indicated by $\overline{\mathcal{M}}$.
We further construct in Fig. 7 a scatter diagram to characterize the degree of ergodicity. The magenta line indicates the ergodic limit $\mathcal{M}/\overline{\mathcal{M}}\sim 1$}. The arrow on top of the scatter diagram indicates that there is a trend towards ‘better’ ergodicity in the increasing $u$ direction. The inspected region of the diagram is where $\overline{\mathcal{M}}$ is large (say more than 20). This is the region where a semiclassical perspective of ergodicity is meaningful. The small value of the $\mathcal{M}/\overline{\mathcal{M}}$ ratio in the MI region is misleading because it is outside of this semiclassical region. Following a more detailed inspection in [18], the reason for getting ‘better’ ergodicity for large $u$ is the shrinking of $\overline{\mathcal{M}}$ (accessible phase-space), while throughout the $\mathcal{M}$ of individual eigenstates is small and much less affected.